Calculate the percent ionization of hypobromous acid (HBrO) in solutions of each of the following concentrations (Ka = 2.5e-09.)

(a) 0.271 M %

(b) 0.388 M %

(c) 0.794 M %
Answer #1

a)

HBrO dissociates as:

HBrO -----> H+ + BrO-

0.271 0 0

0.271-x x x

Ka = [H+][BrO-]/[HBrO]

Ka = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((2.5*10^-9)*0.271) = 2.603*10^-5

since c is much greater than x, our assumption is correct

so, x = 2.603*10^-5 M

% dissociation = (x*100)/c

= 2.603*10^-5*100/0.271

= 9.605*10^-3 %

Answer: 9.60*10^-3 %

b)

HBrO dissociates as:

HBrO -----> H+ + BrO-

0.388 0 0

0.388-x x x

Ka = [H+][BrO-]/[HBrO]

Ka = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((2.5*10^-9)*0.388) = 3.114*10^-5

since c is much greater than x, our assumption is correct

so, x = 3.114*10^-5 M

% dissociation = (x*100)/c

= 3.114*10^-5*100/0.388

= 8.027*10^-3 %

Answer: 8.03*10^-3 %

c)

HBrO dissociates as:

HBrO -----> H+ + BrO-

0.794 0 0

0.794-x x x

Ka = [H+][BrO-]/[HBrO]

Ka = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((2.5*10^-9)*0.794) = 4.455*10^-5

since c is much greater than x, our assumption is correct

so, x = 4.455*10^-5 M

% dissociation = (x*100)/c

= 4.455*10^-5*100/0.794

= 5.611*10^-3 %

Answer: 5.61*10^-3 %